OpenStudy (anonymous):

Mom needs more help please! :Find (-3/7) / (-1 1/14) thank u :):)

OpenStudy (anonymous):

I swear math has changed since I was 13.... lol

OpenStudy (thebatman):

Is this suppose to be in fractional, or decimal form?

OpenStudy (anonymous):

I am assuming fraction. that is all that was in the question

OpenStudy (thebatman):

Just multiply the -3/7 by -14

OpenStudy (superdavesuper):

yup yup: 1 1/14 = 14/14 + 1/14 so that becomes 15/14

OpenStudy (anonymous):

is that the same as 2/5?

OpenStudy (thebatman):

No.

OpenStudy (anonymous):

possible answers I was just given:

OpenStudy (thebatman):

To be honest, again, I don't think I'm reading this question right.

OpenStudy (anonymous):

2/5 or

OpenStudy (superdavesuper):

I will try to write it out...

OpenStudy (superdavesuper):

(-3/7)/(-1 1/14) =(-3/7)/(-15/14) =(-3/7)/((-3x5)/7x2)) =1/5/2 because -3/-3=7/7=1 =2/5

OpenStudy (anonymous):

wow... dave, did u get my message? i was guessing that the answer was 2/5

OpenStudy (anonymous):

it looked right

Parth (parthkohli):

Yeah, dave is correct!

OpenStudy (anonymous):

thanks again dave :)

OpenStudy (superdavesuper):

Welcome mom :)